lucy invests 1320 more at 9% than she invests at 12%. The interest from the two investments is the same. How much did she invest at each rate?

Answer by checkley79(3341)

(X+1320).09=.12X
.09X+118.8=.12X
.09X-.12X=-118.8
-.03X=-118.8
X=-118.8/.03
X=3960 AMOUNT INVESTED @ 12%.
3960+1320=5280 AMOUNT INVESTED @ 9%
PROOF:
(3960+1320).09=.12*3960
5280*.09=475.20
475.2=475.2